OpenStudy (kittiwitti1):

Pre-Calc Questions (Checking because I'm writing an answer key for these problems.) 19 asks for Law of Cosines and area formula. 21 asks for components of a vector with angle of direction given. 46 asks for matrix equations.

OpenStudy (kittiwitti1):

19) Solve the triangle (next comment), then find its area. 21) Find the component form of a vector with magnitude 6 and angle of direction 150°. \[46)~If~D=\left[\begin{matrix}6 & -9 & 7\\ -2 & 1 & 8\end{matrix}\right]~and~F=\left[\begin{matrix}-4 & 3 & 0\\ 6 & 5 & 10\end{matrix}\right],~find~-2D~and~D-F\]

OpenStudy (kittiwitti1):